In the commercial preparation of Hydrogen Chloride, what weight of HCl in grams may be obtained by heating 468 grams of NaCl with H2S04.


2NaCl + H2S04 􏰀 NaSO4 + 2HCl


1
Expert's answer
2022-03-20T06:38:10-0400

"m(HCl)=468\ g(NaCl)\times\frac{1\ mol(NaCl)}{58.44\ g(NaCl)}\times\frac{2\ mol(HCl)}{2\ mol(NaCl)}\times\frac{36.46\ g(HCl)}{1\ mol(HCl)}=292\ g"


Answer: 292 g